THE TENSOR PRODUCT OF POLYNOMIALS

RUTH SCHWINGEL

Abstract. Using Gröbner basis algorithms in MAGMA we find necessary and sufficient conditions for a polynomial of degree 6 over any field to be the tensor product of two polynomials, one of degree 3 and one of degree 2.